an airplane flew for 5 hr with a 20-km/h tailwind and returned in 7 hr against the same wind. find the speed of the plane in still air.
Answer by richwmiller(17219)

5*(x+20)=7*(x-20)
5x+100=7x-140
240=2x
120=x
120 mph in still air